OpenStudy (anonymous):

Find the slope of the line passing through (0, -5) and (-4, 7)

OpenStudy (zzr0ck3r):

(7-(-5))/(-4-0)

OpenStudy (zzr0ck3r):

12/-4 = -3

OpenStudy (lgbasallote):

\[\text{slope} = \frac{y_2 - y_1}{x_2 - x_1}\] \[\implies \frac{7 - (-5)}{-4 - 0} \implies \frac{7 + 5}{-4} \implies \frac{12}{-4} \implies -3\]